# On Viscosity Solutions of Hamilton-Jacobi Equations in the Wasserstein space and the Vanishing Viscosity Limit

### v1

@@ 1.1. Overview on the existing literature
    We begin by briefly discussing the assumptions. The boundedness requirement on the solutions can be relaxed by introducing a suitable growth control; however, we decided not to focus on this technical issue. The main restriction of our result is the further continuity assumption w.r.t. the weak convergence of measures that we briefly discuss here: the space $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is not locally compact, see [AGS, Remark  7.1.9], and this usually constitutes an issue in any optimization procedure, as in the doubling variables argument. However, $\mathcal{P}_{2}(\mathbb{R}^{d})$ is locally compact w.r.t. the weak convergence. We therefore impose further regularity w.r.t. this convergence in order to properly optimize. See the discussion in [bertucci2025doublingvariablestechniqueorder, Section  4.6] where tentative approaches to solve this important problem are also proposed.
    We now heuristically outline the main arguments in the lighter time-independent case, since the core argument is completely analogous to that of the evolutive case. Note that, to derive the vanishing-viscosity limit, we do not rely on possible explicit representations of the solution, as is often done in the control-theoretic setting. Instead, we adopt a purely PDE-based approach, following methods already established in the finite-dimensional case, see the monograph [Bardi1997OptimalCA, Section VI]. The main motivation for this choice is that this approach may also be useful in more general settings, such as the broader metric framework considered in [CKTComp].
    As in the finite dimensional case, we perform a doubling variables argument penalizing the difference between the two solutions:
@@
    where $\mathcal{E}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{+\infty\}$ is the entropy functional relative to the $d$-dimensional Lesbesgue measure $\lambda_{d}$. Thanks to this penalization, we obtain a maximum $(\overline{\mu},\overline{\nu})\in\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})$ (here we use the compactness given by the continuity w.r.t. the weak convergence of the two solutions) at which we can exploit the joint superdifferentiability of $W^{2}_{2}(\cdot,\cdot)$ at $(\overline{\mu},\overline{\nu})$. As usual, this procedure gives elements in the subdifferential of $U^{\varepsilon}-\delta\mathcal{E}$ at $\overline{\mu}$ and in the superdifferential of $U$ at $\overline{\nu}$, which can be used to apply the viscosity solution property to $U^{\varepsilon}$ and $U$. Let us also stress that the maximization gives $\mathcal{I}(\overline{\mu})<\infty$. The viscosity solution property yields
@@
    where the constant $C$ depends only on the local Lipschitz regularity of $H$. The term $C\alpha$ is obtained similarly in the proof of the comparison principle presented in [DS]. The main issue in our case, is the dependence of the constant $K(\varepsilon)$ on the diffusion, which quantitatively compensates the fact that we considered an element of the superdifferential of $U^{\varepsilon}-\delta\mathcal{E}$, and not merely of $U^{\varepsilon}$. Indeed, this constant is expected in general to blowup, see the formal computation in [DS, Section  2.3]. Therefore, we need a bound on $I_{\varepsilon}$ that does not depend on $\delta$. Indeed, applying a naive Cauchy Scwhartz inequality, one can prove that
@@
    Optimizing w.r.t $\delta$ yelds
@@
    as proven in Lemma 4.3, where the regularity assumptions therein are satisfied by the optimizer $\overline{\mu}\in\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})$. This estimate can be regarded as the Wasserstein analogue of the finite-dimensional identity $\varepsilon\Delta_{x}\frac{|y-x|^{2}}{2\alpha}=\frac{d\varepsilon}{\alpha}$ which plays a crucial role in the original proof in [Bardi1997OptimalCA].
@@
    We then conclude the analysis showing that the one side rate of convergence (from above) can be improved from $\sqrt{\varepsilon}$ to $\varepsilon$, provided $U^{\varepsilon}$ is a classical solution such that $\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}\leq C$ uniformly (in the support of $\mu$). In particular, thanks to the analysis on geodesically concave functions, carried out in Section 4, this rate is achieved by $C^{2}$ geodesically $\lambda$-concave functions, at least for $d\geq 2$.
    Finally, the above rates are optimal since they are optimal111w.r.t. the assumptions on the Hamiltonian. See [cirant2025convergenceratesvanishingviscosity, chaintron2025optimalrateconvergencevanishing] for the optimal rate in the convex regime. in the finite dimensional case that embeds in this mean field framework.
@@
In Section 2, we fix the notation, we provide the main assumptions we are going to work with, and recall some important results from the literature that will be useful for our work. Moreover, we investigate some links between various notions of differentials. As a consequence, we will determine one meaningful differential to treat in the viscosity sense both the equations (1.2) and (1.2).
@@ 2.1. The Wasserstein space
is the transport cost from $\mu$ to $\nu$ associated to $\gamma$. For further details, see [OTVillani, Chapter 4, Chapter 6].
@@
$\Gamma({\gamma},\nu):=\{\theta\in\mathcal{P}(\mathbb{R}^{3d}):(\pi_{1},\pi_{2})_{\#}\theta=\gamma,\ (\pi_{1},\pi_{3})_{\#}\theta\in\Gamma({(\pi_{1})}_{\#}\gamma,\nu)\}$
@@
To any $\gamma\in\mathcal{P}(\mathbb{R}^{2d})$ we associate its disintegration w.r.t. the first marginal $\mu:=(\pi_{1})_{\#}\gamma$: i.e. the $\mu$-a.e. uniquely determined measure $\gamma_{x}(dy)\in\mathcal{P}(\mathbb{R}^{d})$ s.t. $\gamma=\gamma_{x}(dy)\mu(dx)$. The disintegration easily generalizes to the second marginal and to every marginal of measures $\theta\in\mathcal{P}(\mathbb{R}^{3d})$. See [AGS, Theorem 5.3.1].
@@
For future use, we mention that for any $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, as a simple consequence of the Cauchy-Schwartz inequality we have the following inequalities
@@
$\leq W_{2}(\mu,\nu)\left((\int_{\mathbb{R}^{d}}|x|^{2}d\mu)^{\frac{1}{2}}+(\int_{\mathbb{R}^{d}}|y|^{2}d\nu(y))^{\frac{1}{2}}\right).$
@@ 2.2. Tangent spaces at a measure
We survey here the definitions of various notions of tangent space at a measure, as described in [AGS, Appendix], the definition of tangent bundle and their links. Thanks to this description, we can unify the notions of viscosity solution in the deterministic case and in the presence of idiosyncratic noise.
@@
The set $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ is metrized by the distance $W_{\mu}\colon\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})\times\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})\to[0,\infty)$ defined in [AGS, Equation  (12.4.10)], as
@@
Note that the convergence induced by $W_{\mu}$ on $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ coincides with that induced by the natural extension of $W_{2}$ on $\mathcal{P}_{2}(\mathbb{R}^{2d})$ restricted to $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$. [^2] To see this: Let us consider the natural extension of $W_{2}$ on $\mathcal{P}_{2}(\mathbb{R}^{2d})$ defined as $W^{2}_{2}(\gamma_{1},\gamma_{2})=\min_{\theta\in\Gamma(\gamma_{1},\gamma_{2})}\int_{\mathbb{R}^{4d}}(|x_{2}-x_{1}|^{2}+|z_{1}-z_{2}|^{2})d\theta(x_{1},z_{1};x_{2},z_{2})$ for $\gamma_{1},\gamma_{2}\in\mathcal{P}_{2}(\mathbb{R}^{2d})$. Here $\Gamma(\gamma_{1},\gamma_{2})=\big\{\sigma\in\mathcal{P}(\mathbb{R}^{2d}\times\mathbb{R}^{2d}):(\pi_{i})_{\#}\sigma=\gamma_{i},i={1,2}\big\},$ where $\pi_{i}:\mathbb{R}^{2d}\times\mathbb{R}^{2d}\to\mathbb{R}^{2d},i=1,2$ are the natural projections to the first and second factor, respectively. We note that $\int_{\mathbb{R}^{4d}}(|x_{2}-x_{1}|^{2}+|z_{1}-z_{2}|^{2})d\theta(x_{1},z_{1};x_{2},z_{2})\geq\int_{\mathbb{R}^{4d}}|z_{1}-z_{2}|^{2}d\theta(x_{1},z_{1};x_{2},z_{2})$ $\forall\theta\in\Gamma(\gamma_{1},\gamma_{2})\subset\mathcal{P}(\mathbb{R}^{4d})$, and the equality holds iff $x_{1}=x_{2},\,\theta$-a.e., i.e. $\gamma_{1},\gamma_{2}$ shares the same first marginal and $(\pi_{1},\pi_{3})_{\#}\theta=(\text{Id}\times\mathrm{Id})_{\#}\mu$. In particular, $\theta$ can be identified as a plan in $\mathcal{P}(\mathbb{R}^{3d})$ belonging to $\Gamma_{\mu}(\gamma_{1},\gamma_{2})$, and $W_{\mu}$ coincides with the restriction of the $W_{2}$ to $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$.
@@
The (Kantorovich) geometric tangent space at $\mu$ is defined as:
@@
In $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$, it is naturally defined an exponential map
@@
We remark that, as in [GigliPhd, Definition  4.18], we prefer to define the exponential on the whole $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ rather than on its closed subset $\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ for reasons that will become clear later in our discussion about differentials. Moreover, let us note that $\exp_{\mu}$ is surjective, i.e., more precisely, for all $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ there exists an optimal velocity $\gamma\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ s.t. $\exp_{\mu}(\gamma)=\nu$.
Following [GigliPhd, Definition  4.13], given $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, we can also introduce a notion of “scalar product” on $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ as
@@
By [GigliPhd, Proposition  4.21], we have that, for $\gamma_{1},\gamma_{2},\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$,
@@
In other words, the scalar product is Lipschitz continuous w.r.t. $W_{\mu}$. We also remark, see [GigliPhd, Remark  4.28], that in general $\max_{\theta\in\Gamma_{\mu,\nu}(\gamma,\xi)}\big<\gamma,\xi\big>_{\theta,\mu}>\min_{\theta\in\Gamma_{\mu,\nu}(\gamma,\xi)}\big<\gamma,\xi\big>_{\theta,\mu}.$ However, the equality holds if either $\gamma$ or $\xi$ is induced by a map. Lastly, it is defined a multiplication operation, see in [GigliPhd, Proposition  4.22 &\& Proposition  4.24]: given $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
@@
If $\gamma\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$, it is immediate to see that $\lambda\cdot\gamma\in\mathrm{Tan}_{\mu}^{K}{\mathcal{P}}_{2}(\mathbb{R}^{d})$, for every $\lambda\geq 0$. This invariance is still true for $\lambda<0$, see [GigliPhd, Proposition  4.29]. To lighten the notation we will often write $\gamma$ for $1\cdot\gamma$. We also remark that the product $\big<\cdot,\cdot\big>_{\mu}$ enjoys a one side Cauchy-Schwarz inequality: Indeed, for every $\gamma^{1},\gamma^{2}\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
$\big<\gamma^{1},\gamma^{2}\big>_{\mu}$ $=\max_{\theta\in\Gamma_{\mu}(\gamma^{1},\gamma^{2})}\int\big<w_{1},w_{2}\big>d\theta(x,w_{1},w_{2})$
$\leq\max_{\theta\in\Gamma_{\mu}(\gamma^{1},\gamma^{2})}\Big(\int|w_{1}|^{2}d\theta(x,w_{1},w_{2})\Big)^{\frac{1}{2}}\Big(\int|w_{2}|^{2}d\theta(x,w_{1},w_{2})\Big)^{\frac{1}{2}}$
$=\Big(\int|w_{1}|^{2}d\gamma^{1}(x,w_{1})\Big)^{\frac{1}{2}}\Big(\int|w_{2}|^{2}d\gamma^{2}(x,w_{2})\Big)^{\frac{1}{2}}$
@@
Moreover, the equality holds iff there exists $\theta\in\Gamma_{\mu}(\gamma^{1},\gamma^{2})$ optimizer for $\big<\cdot,\cdot\big>_{\mu}$ s.t. $\big<w_{1},w_{2}\big>=|w_{1}||w_{2}|$ holds $\theta$-a.e., i.e. $\gamma^{1}=\lambda\gamma^{2}$, with $\lambda\geq 0$.
As in [GigliPhd, Proposition  4.30], we define a projection operator on $\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{2d})$:
@@
that at each element $\gamma$ associates the unique plan $\text{P}(\gamma)\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ that minimizes the distance between $\gamma$ and the geometric tangent space $\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$. In particular, see [GigliPhd, Corollary  4.34],
@@
In order to understand the link between the two different notions of tangent space, let us consider the barycentric projection (see [AGS, Definition  5.4.2.]). For a fixed $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the barycentric projection map is defined as
@@
$\overline{\gamma}(x)=\int_{\mathbb{R}^{d}}y\,d\gamma_{x}(y).$
Note that Jensen inequality ensures the wellposedness of the above map, indeed $\overline{\gamma}\in L^{2}_{\mu}(\mathbb{R}^{d})$. To be more general, we can define the projection map [^3] In a probabilistic framework, the barycentric projection of the law of a random vector $(X,Y)$ corresponds to the law of the conditional expectation $\mathbb{E}[Y|X]$. on the whole $\mathcal{P}_{2}(\mathbb{R}^{2d})$, as
$\text{Pr}(\gamma):=\text{Pr}_{(\pi_{1})_{\#}\gamma}(\gamma),\quad\forall\gamma\in\mathcal{P}_{2}(\mathbb{R}^{2d}).$
It turns out that, for any $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the Monge tangent space can be seen as the image of the restriction of the barycentric projection to the geometric tangent space. Moreover, it is not difficult to show that Pr is onto and a 11-contraction (see [AGS, Definition 12.4.3], and the subsequent discussion). Namely
@@
When $\mu$ is absolutely continuous, we can be more precise: this projection is a one-to-one isometry w.r.t. the distances previously introduced, as shown in [AGS, Theorem  12.4.4]. Thus
@@
Let us note that, using the characterization of optimal maps for absolutely continuous measures, we also have the following alternative description for the tangent space (see [AGS, Definition  8.4.1 &\& Theorem  8.5.1]):

### v2

@@ 1.1. Overview on the existing literature
    We begin by briefly discussing the assumptions. The boundedness requirement on the solutions can be relaxed by introducing a suitable growth control; however, we decided not to focus on this technical issue. The main restriction of our result is the further continuity assumption w.r.t. the weak convergence of measures that we briefly discuss here: the space $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is not locally compact, see [2, Remark  7.1.9], and this usually constitutes an issue in any optimization procedure, as in the doubling variables argument. However, $\mathcal{P}_{2}(\mathbb{R}^{d})$ is locally compact w.r.t. the weak convergence, recalled in Subsection 2.6. We therefore impose further regularity w.r.t. this convergence in order to properly optimize. See also the discussion in [8, Section  4.6] where tentative approaches to solve this important problem are also proposed.
    We now heuristically outline the main arguments in the lighter time-independent case, since the core argument is completely analogous to that of the evolutive case. Note that, to derive the vanishing-viscosity limit, we do not rely on possible explicit representations of the solution, as is often done in the control-theoretic setting. Instead, we adopt a purely PDE-based approach, following methods already established in the finite-dimensional case, see the monograph [5, Section VI] or [23, Theorem 5.1]. The main motivation for this choice is that this approach may also be useful in more general settings, such as the broader metric framework considered in [21].
    Then the strategy closely follows the finite dimensional case. We perform a doubling variables argument penalizing the difference between the two solutions:
@@
    where $\mathcal{E}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{+\infty\}$ is the entropy functional relative to the $d$-dimensional Lesbesgue measure $\lambda_{d}$. Thanks to this penalization, we obtain a maximum $(\overline{\mu},\overline{\nu})\in\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})$ at which we can exploit the joint superdifferentiability of $W^{2}_{2}(\cdot,\cdot)$ at $(\overline{\mu},\overline{\nu})$ (here we use the compactness ensured by the continuity w.r.t. the weak convergence of the two solutions, indeed the $W_{1}$-Lipschitz continuity of $U^{\varepsilon}$ implies its continuity w.r.t. weak convergence). As usual, this procedure gives elements in the subdifferential of $U^{\varepsilon}-\delta\mathcal{E}$ at $\overline{\mu}$ and in the superdifferential of $U$ at $\overline{\nu}$, which can be used to apply the viscosity solution property to $U^{\varepsilon}$ and $U$. Let us also stress that the maximization gives $\mathcal{I}(\overline{\mu})<\infty$. We then apply the viscosity solution property which yields
@@
    where the constant $C$ depends only on the local Lipschitz regularity of $H$. The term $C\alpha$ is obtained similarly in the proof of the comparison principle presented in [30]. The main issue in our case, is the dependence of the constant $K(\varepsilon)$ on the diffusion, which quantitatively compensates the fact that we considered an element of the superdifferential of $U^{\varepsilon}-\delta\mathcal{E}$, and not merely of $U^{\varepsilon}$. Indeed, this constant is expected in general to blowup, see the formal computation in [30, Section  2.3, Eq. (2.23)]. Therefore, we need a bound on $I_{\varepsilon}$ that does not depend on $\delta$. Indeed, applying a naive Cauchy Schwarz inequality, one can prove that
@@
    Optimizing w.r.t $\delta$ yields
@@
    as proven in Lemma 4.4, where the regularity assumptions therein are satisfied by the optimizer $\overline{\mu}\in\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})$. This estimate can be regarded as the Wasserstein analogue of the finite-dimensional identity $\varepsilon\Delta_{x}\frac{|y-x|^{2}}{2\alpha}=\frac{d\varepsilon}{\alpha}$ which plays a crucial role in the finite dimensional proof in [5]. The inequality can be interpreted as a manifestation of the positive curvature of the space $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$, as observed in Remark (4.5).
@@
    We then conclude the analysis showing that the one sided rate of convergence (from above) can be improved from $\sqrt{\varepsilon}$ to $\varepsilon$, provided $U^{\varepsilon}$ is a classical solution of (1.2) such that $\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(t,\mu,x)\leq C$ uniformly. In particular, thanks to the analysis on geodesically concave functions, carried out in Section 4, this rate is achieved by $C^{2}$ geodesically $\Lambda$-concave functions, at least for $d\geq 2$.
    Finally, the above rates are optimal since they are optimal111w.r.t. the assumptions on the Hamiltonian. See [19, 18] for the optimal rate in the convex regime. in the finite dimensional case that embeds in this mean field framework.
@@
In Section 2, we fix the notation, we provide the main assumptions we are going to work with, and recall some important results from the literature that will be useful for our work. Moreover, we investigate some links between various notions of differentials. As a consequence, we will determine one meaningful differential to treat both the equations (1.2) and (1.2), in the viscosity sense.
@@ 2.1. The Wasserstein Space
is the transport cost from $\mu$ to $\nu$ associated to $\gamma$. For further details, see [50, Chapter 4, Chapter 6].
@@
$\Gamma({\gamma},\nu):=\{\theta\in\mathcal{P}(\mathbb{R}^{3d}):(\pi_{1},\pi_{2})_{\#}\theta=\gamma,\ (\pi_{1},\pi_{3})_{\#}\theta\in\Gamma({(\pi_{1})}_{\#}\gamma,\nu)\},$
@@
To any $\gamma\in\mathcal{P}(\mathbb{R}^{2d})$ we associate its disintegration w.r.t. the first marginal $\mu:=(\pi_{1})_{\#}\gamma$: i.e. the $\mu$-a.e. uniquely determined measure $\gamma_{x}(dy)\in\mathcal{P}(\mathbb{R}^{d})$ s.t. $\gamma=\gamma_{x}(dy)\mu(dx)$. The disintegration easily generalizes to the second marginal and to every marginal of measures $\theta\in\mathcal{P}(\mathbb{R}^{3d})$. See [2, Theorem 5.3.1].
@@
For future use, we mention that for any $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, as a simple consequence of the Cauchy Schwarz inequality we have the following inequalities
@@
$\leq W_{2}(\mu,\nu)\left((\int_{\mathbb{R}^{d}}|x|^{2}d\mu(x))^{\frac{1}{2}}+(\int_{\mathbb{R}^{d}}|y|^{2}d\nu(y))^{\frac{1}{2}}\right).$
Note that the previous notions can be easily generalized to $\mathcal{M}_{p}(\cdot),\mathcal{P}_{p}(\mathbb{R}^{d}),W_{p}(\cdot,\cdot),\mathrm{Lip}(\cdot;W_{p})$ for $p\geq 1$.
@@ 2.2. Tangent spaces at a measure
We survey here the definitions of various notions of tangent space at a measure, as described in [2, Appendix]. Thanks to this description, we can unify the notions of viscosity solution in the first order case and in the semilinear case with idiosyncratic noise.
@@
The set $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ is metrized by the distance $W_{\mu}\colon\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})\times\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})\to[0,\infty)$ defined in [2, Equation  (12.4.10)], as
@@
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. The (Kantorovich) geometric tangent space at $\mu$ is defined as:
@@
In $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$, it is naturally defined an exponential map:
@@
We remark that, as in [37, Definition  4.18], we prefer to define the exponential on the whole $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ rather than on its closed subset $\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ for reasons that will become clear later in our discussion about differentials. Moreover, let us note that $\exp_{\mu}$ is surjective, i.e., more precisely, for all $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ there exists an optimal velocity $\gamma\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ s.t. $\exp_{\mu}(\gamma)=\nu$.
Following [37, Definition  4.13], given $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, we can also introduce a notion of “scalar product” on $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ as
@@
By [37, Proposition  4.21], we have that, for $\gamma_{1},\gamma_{2},\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$,
@@
In other words, the scalar product is Lipschitz continuous w.r.t. $W_{\mu}$. We also remark, see [37, Remark  4.28], that in general
(2.7) $\max_{\theta\in\Gamma_{\mu,\nu}(\gamma,\xi)}\big<\gamma,\xi\big>_{\theta,\mu}\geq\min_{\theta\in\Gamma_{\mu,\nu}(\gamma,\xi)}\big<\gamma,\xi\big>_{\theta,\mu},$
with the inequality possibly strict. However, the equality holds if either $\gamma$ or $\xi$ is induced by a map. Lastly, a multiplication operation can be defined on $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$, see [37, Proposition  4.22 &\& Proposition  4.24]: given $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
@@
If $\gamma\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$, it is immediate to see that $\lambda\cdot\gamma\in\mathrm{Tan}_{\mu}^{K}{\mathcal{P}}_{2}(\mathbb{R}^{d})$, for every $\lambda\geq 0$. This invariance is still true for $\lambda<0$, see [37, Proposition  4.29]. To lighten the notation we will often write $\gamma$ for $1\cdot\gamma$. We also remark that the product $\big<\cdot,\cdot\big>_{\mu}$ enjoys a one sided Cauchy Schwarz inequality: Indeed, for every $\gamma^{1},\gamma^{2}\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
$\big<\gamma^{1},\gamma^{2}\big>_{\mu}$ $=\max_{\theta\in\Gamma_{\mu}(\gamma^{1},\gamma^{2})}\int\big<y,z\big>d\theta(x,y,z)$
$\leq\max_{\theta\in\Gamma_{\mu}(\gamma^{1},\gamma^{2})}\Big(\int|y|^{2}d\theta(x,y,z)\Big)^{\frac{1}{2}}\Big(\int|z|^{2}d\theta(x,y,z)\Big)^{\frac{1}{2}}$
$=\Big(\int|y|^{2}d\gamma^{1}(x,y)\Big)^{\frac{1}{2}}\Big(\int|z|^{2}d\gamma^{2}(x,z)\Big)^{\frac{1}{2}}$
@@
Moreover, the equality holds iff there exists $\theta\in\Gamma_{\mu}(\gamma^{1},\gamma^{2})$ optimizer for $\big<\cdot,\cdot\big>_{\mu}$ s.t. $\big<y,z\big>=|y||z|$ holds $\theta$-a.e., i.e. $\gamma^{1}=\lambda\cdot\gamma^{2}$, with $\lambda\geq 0$.
As in [37, Proposition  4.30], we define a projection operator on $\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{2d})$ as:
@@
that at each element $\gamma$ associates the unique plan $\text{P}(\gamma)\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ that minimizes the distance between $\gamma$ and the geometric tangent space $\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$. In particular, see [37, Corollary  4.34],
@@
In order to understand the link between the two different notions of tangent space, let us consider the barycentric projection (see [2, Definition  5.4.2.]). For a fixed $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the barycentric projection map is defined as
@@
$\overline{\gamma}(x):=\int_{\mathbb{R}^{d}}y\,d\gamma_{x}(y).$
Note that Jensen inequality ensures the wellposedness of the above map, indeed $\overline{\gamma}\in L^{2}_{\mu}(\mathbb{R}^{d})$. To be more general, we define the projection map [^2] In a probabilistic framework, the barycentric projection of the law of a random vector $(X,Y)$ corresponds to the law of the conditional expectation $\mathbb{E}[Y|X]$. on the whole $\mathcal{P}_{2}(\mathbb{R}^{2d})$, as
(2.9) $\text{Pr}(\gamma):=\text{Pr}_{(\pi_{1})_{\#}\gamma}(\gamma),\quad\forall\gamma\in\mathcal{P}_{2}(\mathbb{R}^{2d}).$
It turns out that, for any $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the Monge tangent space can be seen as the image of the restriction of the barycentric projection to the geometric tangent space. Moreover, it is not difficult to show that Pr is onto and a 11-contraction (see [2, Definition 12.4.3], and the subsequent discussion). Namely
@@
When $\mu$ is absolutely continuous, we can be more precise: this projection is a one-to-one isometry w.r.t. the distances previously introduced, as shown in [2, Theorem  12.4.4]. Thus
@@
Let us note that, using the characterization of optimal maps for absolutely continuous measures, we also have the following alternative description for the tangent space (see [2, Definition  8.4.1 &\& Theorem  8.5.1]):